2 p 7 0 The angles and the mass M given Find the tesion in each on the pic lure 90e af the chains L, M and N 30 Kg OVE ER
Answer #1

here,

m = 30 kg

the tesnion in the cable N , Tn = m * g

Tn = 30 * 9.81 N = 294.3 N

and

equating the forces vertically

Tm * sin(70) = Tn

Tm * sin(70) = 294.3 N

Tm = 313.2 N

equating the forces horizontally

Tl = Tm * cos(70)

Tl = 313.2 * cos(70) N = 107.1 N